OpenStudy (anonymous):

A theatre can seat 280 people. The number of rows is 6 less than the number of seats in each row. How many rows of seats are there?????

OpenStudy (anonymous):

x(x-6) =280 x^2 -6X -280= 0 (x-20)(x+14)=0 x=20 or -14 (rej) number of rows = 20-6=14

OpenStudy (anonymous):

thankyouthankyou

OpenStudy (anonymous):

you are welcome